OpenStudy (anonymous):

JASON IS A SALES PERSON IN THE ELECTRONICS DEPARTMENT OF A STORE. HE IS 30% SHORT OF HIS GOAL FOR THE WEEK. HIS GOAL IS TO EARN $500.FOR EACH CD HE SELLS, HE EARNS $25. HOW MANY CD PLAYERS DOES HE NEED TO SELL IN ORDER TO MEET HIS GOAL? a minute ago

OpenStudy (anonymous):

He needs to earn 30% of 500 more , 30% of 500 = 3/10 * 500 = $150 no. of CDs he needs to sell = 150/25 = 6

OpenStudy (anonymous):

THANKS